OpenStudy (anonymous):

Sqrt 25x divided by sqrt 5y. Simplify. Show your work.

OpenStudy (mertsj):

\[\frac{\sqrt{25x}}{\sqrt{5y}}=\frac{\sqrt{25}}{\sqrt{5}}\times\frac{\sqrt{x}}{\sqrt{y}}=\frac{\sqrt{5x}}{\sqrt{y}}\times\frac{\sqrt{y}}{\sqrt{y}}=\frac{\sqrt{5xy}}{y}\]